Which of the following is a likely advantage in activity-based costing (ABC)?

  1. An increase in the cost of accumulating, tracking, and assigning costs to products and services.   
  2. An increase in the ability to trace indirect costs to specific activities.        
  3. A decrease in the need to maintain accurate cost driver information.      
  4. A decrease in the need to maintain accurate overhead cost information.

Which of the following is true regarding activity-based costing (ABC)?

  1. Companies with a high proportion of non-unit-level costs should rarely use ABC.
  2. High measurement costs associated with ABC may deter some companies from its use.  
  3. The benefits of using ABC always outweigh its limitations.             
  4. The lower the proportion of overhead costs as compared to direct materials and direct labor costs, the more likely a company will benefit by using ABC.       

Cornell Products has the following information available for 2012:

Direct materials                                                                $1.00 per unit

Direct labor                                                                         $2.00 per unit

Variable manufacturing overhead                            $1.50 per unit

Variable selling and administrative costs                $0.50 per unit

Fixed manufacturing overhead                                  $30,000

Fixed selling and administrative costs                      $25,000

      During 2012, Cornell produced 6,000 units out of which 5,400 units were sold for $20 each.

      Refer to the information provided for Cornell Products. What is the net operating income under
      absorption costing?

  1. $26,000                
  2. $35,000                
  3. $23,000                
  4. $29,000                

Hillsborough Street Manufacturing Inc.

Hillsborough Street Manufacturing Inc. incurred the following costs in 2011:

Direct materials used                     $51,000

Direct labor costs                             45,000

Factory rent and utilities               18,000

Factory equipment depreciation    7,500

Marketing expenses                           8,000

Administrative expenses              10,000

      45,000 units were produced during the year out of which 38,000 units were sold for $10 each. There
      was no beginning or ending raw materials or work in process inventory.

      Refer to the Hillsborough Street Manufacturing Inc. information above. What is the net operating
      income for the year? (Ignore taxes)

                a. $222,500         

                b. $240,500        

                c. $244,200         

                d. $259,400        

Hillsborough Street Manufacturing Inc.

Hillsborough Street Manufacturing Inc. incurred the following costs in 2011:

Direct materials used                     $51,000

Direct labor costs                             45,000

Factory rent and utilities               18,000

Factory equipment depreciation                   7,500

Marketing expenses                         8,000

Administrative expenses              10,000

      45,000 units were produced during the year out of which 38,000 units were sold for $10 each. There
      was no beginning or ending raw materials or work in process inventory.

      Refer to the Hillsborough Street Manufacturing Inc. information above. What is cost of goods sold for
      the year?

  1. $121,500             
  2. $117,800             
  3. $139,500             
  4. $102,600             
Answer #1

Answer 1:

Correct answer is:

b. An increase in the ability to trace indirect costs to specific activities

Explanation:

ABC defines activities and activity cost pools. It enhances ability to trace indirect costs to specific activities.

As such option b is correct.

Option a is incorrect since "An increase in the cost of accumulating, tracking, and assigning costs to products and services" is a disadvantage of ABC.

Options c and d are incorrect since ABC increases the need to maintain accurate cost driver information and need to maintain accurate overhead cost information.

Answer 2:

Correct answer is:

b. High measurement costs associated with ABC may deter some companies from its use.

Explanation:

Usage of ABC system needs definition of activities, activity cost pools, cost drivers, cost driver rates and measurement of the same. For smaller companies and for companies with low proportion of overheads may deter them from its use due to high measurement costs associated with it. Since in these cases benefits may not outweigh costs associated with it.

As such option b is correct.

Option a is incorrect since companies with a high proportion of non-unit-level costs should use ABC.

Option c is incorrect since in case of smaller companies benefits of using ABC may not outweigh its limitations.

Option d is incorrect since the higher the proportion of overhead costs as compared to direct materials and direct labor costs, the more likely a company will benefit by using ABC.

Answer 3:

Correct answer is:

d. $259,400

Explanation:

Manufacturing cost = Direct materials used + Direct labor costs + Factory overheads = 51000 + 45000 + 18000 + 7500 = $121,500

Finished goods (FG) = (45000 - 38000) * 121500 / 45000 = $18,900

Cost of goods sold (COGS) = Beginning FG inventory + Manufacturing cost - Ending FG inventory = 0 + 121500 - 18900 = $102,600

Net operating income = Sales - COGS - Marketing expenses - Administrative expenses

= 38000 * $10 - 102600 - 8000 - 10000

= $259,400

As such option d is correct and other options a, b and c are incorrect.

Answer 4:

Correct answer is:

d. $102,600   

Explanation:

Manufacturing cost = Direct materials used + Direct labor costs + Factory overheads = 51000 + 45000 + 18000 + 7500 = $121,500

Finished goods (FG) = (45000 - 38000) * 121500 / 45000 = $18,900

Cost of goods sold (COGS) = Beginning FG inventory + Manufacturing cost - Ending FG inventory = 0 + 121500 - 18900 = $102,600

As such option d is correct and other options a, b and c are incorrect.